# Existence, uniqueness, and regularity of

### v1

@@ 3.2 Stability Theorem
$\begin{cases}u^{n_{j}}_{t}+{F}\left(t,x,u^{n_{j}},u_{x}^{n_{j}},u_{xx}^{n_{j}}\right)\leq-{h^{1}}(t,x)\text{ in }\mathcal{V}_{n_{j}}\\ u^{n_{j}}=g^{1}\text{ on }\partial\mathcal{V}_{n_{j}}.\end{cases}$
But, by definition of ${h^{1}}$, $g^{1}\in W^{1,2,p}\left(\mathcal{Y}\right)$ is an $L^{p}$-solution of
$\begin{cases}g^{1}_{t}+{F}\left(t,x,g^{1},g^{1}_{x},g^{1}_{xx}\right)=-{h^{1}}(t,x)\text{ in }\mathcal{V}_{n_{j}}\\ g^{1}=g^{1}\text{ on }\partial\mathcal{V}_{n_{j}}.\end{cases}$
Therefore, by our comparison principle (Proposition 1), $g^{1}\leq u^{n_{j}}$ on $\bar{\mathcal{V}}_{n_{j}}$: a contradiction. So, for all $j\in\mathbb{N}$, $\mathcal{V}_{n_{j}}=\emptyset$, i.e., $u^{n_{j}}\geq g^{1}$ on $\bar{\mathcal{Y}}$. Taking limits, as $\left(u^{n_{j}}\right)_{j\in\mathbb{N}}$ converges uniformly to $u$ on $\bar{\mathcal{Y}}$, $u\geq g^{1}$ on $\mathcal{Y}$.
So far, we have shown that $u=b$ on $\partial\mathcal{Y}$ and that
$\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g^{1}-u\right\}\leq 0\text{ a.e. in }\mathcal{Y},$
i.e., $u$ is an $L^{p}$-supersolution of (1), with $u=b$ on $\partial\mathcal{Y}$. To conclude, there remains to show that
$\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g^{1}-u\right\}\geq 0\text{ a.e. in }\mathcal{Y}.$
@@
By definition of $\Phi_{\epsilon}$, note that $\Phi_{\epsilon_{n_{j}}}\left(u^{{n_{j}}}(t,x)-g(t,x)\right)\to 1$ pointwise a.e. as $j\to\infty$ on $\mathcal{U}$. Therefore, for all $(t_{0},x_{0})\in\mathcal{U}$ and $D>0$ such that $C_{D}(t_{0},x_{0})\subset\mathcal{U}$, by Theorem 2, $u$ is an $L^{p}$-subsolution of
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)=0\text{ in }C_{r}(t_{0},x_{0}).$
@@
$\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g^{1}-u\right\}\geq 0\text{ a.e. in }\mathcal{U}.$
So,
$\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g^{1}-u\right\}\geq 0\text{ a.e. in }\mathcal{Y}.$
and $u$ is an $L^{p}$-solution of (1) with $I=1$. This concludes the base case of our induction argument.
Next, suppose that $P(I)$ holds for some $I\in\mathbb{N}$. We show that it also holds for $I+1$. Let $g^{I+1}\in W^{1,2,p}\left(\mathcal{Y}\right)$ with $g^{I+1}\leq b$ on $\partial\mathcal{Y}$. Define $h^{I+1}=-g^{I+1}_{t}-F\left(t,x,g^{I+1},g^{I+1}_{x},g^{I+1}_{xx}\right)$; and consider the following sequence of auxiliary problems. For all $n\in\mathbb{N}$,
$\begin{cases}&\max\bigg{\{}u^{n}_{t}+F\left(t,x,u^{n},u^{n}_{x},u^{n}_{xx}\right)-h^{I+1}(t,x)^{+}\Phi_{\epsilon_{n}}\left(u^{{n}}-g^{I+1}\right)+h^{I+1}(t,x)^{+},\\ &\qquad\qquad\qquad\qquad\qquad\qquad\underset{i\in\{1,\dots,I\}}{\max}\,g^{i}-u^{n}\bigg{\}}=0\text{ in }\mathcal{Y},\\ &u^{n}=b\text{ on }\partial\mathcal{Y}.\end{cases}$ (DI+1)
@@
Proceeding exactly as in the case $I=1$, we see that the sequence $\left(u^{n}\right)_{n\in\mathbb{N}}$ has a subsequence that converges weakly in $W^{1,2,p}\left(\mathcal{Y}\right)$ and strongly in $C^{0}\left(\bar{\mathcal{Y}}\right)$ to some $u\in W^{1,2,p}\left(\mathcal{Y}\right)$. Moreover, following the same steps as above, one can show that $u$ is an $L^{p}$-solution of (1) with $I+1$ obstacles.
To conclude, there only remains to show that the bound (6) on the $W^{1,2,p}$-norm of $u$ holds. Note that $u_{t}=g_{t}$, $u_{x}=g_{x}$, and $u_{xx}=g_{xx}$ a.e. on $\left\{u=g\right\}$. This follows from the proof of Corollary 3.1.2.1 in Evans, (2018), using Proposition A.1 in Crandall et al., (1998) instead of Rademacher’s Theorem. So, $u\in{W}^{1,2,p}\left(\mathcal{Y}\right)$ solves
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)=\mathbbm{1}_{\left\{u>\underset{i=1,\dots,I+1}{\max}\,g^{i}(t,x)\right\}}h^{i(t,x)}(t,x)-h^{i(t,x)}(t,x),$ (8)
where $h^{i}(t,x)=-g^{i}_{t}-F\left(t,x,g^{i},g^{i}_{x},g^{i}_{xx}\right)$ and $i(t,x)$ is a measurable selection from $\arg\underset{i\in\{1,\dots I+1\}}{\max}\,g^{i}(t,x)$. The result then follows from Theorem 15.1.3 in Krylov, (2018) (using Lemma 12.1.9 in Krylov, (2018) to bound $\left\|u\right\|_{L^{\infty}\left(\mathcal{Y}\right)}$ and Theorem 9.8.1 in Krylov, (2018) to obtain the modulus of continuity of $u$).
@@
Let $p\in(d+2,\infty]$ and suppose that Assumptions 1’, and 27 hold and that $\left|A\right|<\infty$. If $u$ is an $L^{p}$-solution of (1), then, for all compact subset $\mathcal{Y}^{\prime}$ of $\mathcal{Y}$,
@@ 5 Proof of Theorem 1
We first prove the result for the special case of smooth domains.
@@
Let $p\in(d+2,\infty]$ and suppose that Assumptions 1’, 2’, and 37 hold. Then (1) has an $L^{p}$-solution $u$.
@@
5.1 Proof of Lemma 3
Building on Lemma 2, we We approximate the obstacle $g=\underset{a\in A}{\sup g^{a}}$ by a sequence of obstacles that satisfy the Assumptions of Lemma 2. We then invoke the stability result for obstacle problems derived in Section 3.2 (Theorem 2) to conclude.
@@ 5.1 Proof of Lemma 3
$\begin{cases}\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,\underset{a\in A^{0,n}}{\sup}\,g^{a}-u\right\}=0\text{ in }\mathcal{Y},\\ u=b\text{ on }\partial\mathcal{Y}.\end{cases}$
@@
Since $W^{1,2,p}(\mathcal{Y})$ is separable and reflexive, its closed bounded subsets are weakly sequentially compact by Theorem 1.32 in Demengel et al., (2012). Moreover $W^{1,2,p}\left(\mathcal{Y}\right)$ is compactly embedded in $\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$ by Rellich-Kondrachov’s theorem (Theorem 2.84 in Demengel et al., (2012)). So, there exists a function $u\in W^{1,2,p}\left(\mathcal{Y}\right)\cap\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$ and a subsequence $\left(u^{n_{j}}\right)_{j\in\mathbb{N}}\subseteq\left(u^{n}\right)_{n\in\mathbb{N}}$ such that
@@
To conclude, there only remains to show that $u$ is the solution to (1), which follows from Theorem 2. $\blacksquare$
From the proof of Lemma 3 and Corollary 2, we obtain the following interior $W^{1,2,p}$-estimates and $L^{\infty}$ bound.
@@
Let $p\in(d+2,\infty]$ and suppose that Assumptions 1’, and 27 hold. If $u$ is an $L^{p}$-solution of (1), then, for all compact subset $\mathcal{Y}^{\prime}$ of $\mathcal{Y}$,
@@ 5.2 Proof of Theorem 1
Lemma 3 guarantees that the obstacle problem has an $L^{p}$-solution when $\mathcal{X}$ and $b$ satisfy Assumption 1’ and 2’. To generalize the result to the weaker Assumptions 1 and 2, we study a sequence of equations, each satisfying the Assumptions of Lemma 3, that converges to the equation (1). In particular, we approximate (i) $\mathcal{X}$ by a sequence of smooth domains whose cone parameters are uniformly controlled, and (ii) $b$ by a sequence of equicontinuous functions in $W^{1,2,p}\left(\mathcal{Y}\right)$. The $L^{p}$ solutions of the equations in the approximating sequence form an equicontinuous (Lemma 4) and weakly compact (in $W^{1,2,p}_{loc}\left(\mathcal{Y}\right)$) family. Arzela-Ascoli’s theorem then guarantees that a subsequence converges uniformly on the compact subset of $\bar{\mathcal{Y}}$ to some function in $W^{1,2,p}_{loc}\left(\mathcal{Y}\right)\cap\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$. Finally, we invoke the stability result for obstacle problems (Theorem 2) derived in Section 3.2 to conclude.
@@
Let $p\in(d+2,\infty]$ and suppose that Assumptions 1’, 2’, and 37 hold. Let $u$ be an $L^{p}$-solution of (1). For all $D>0$, there exists a family of modulus of continuity $\bar{\omega}_{D}$ such that, for all $(t,x),(t^{\prime},x^{\prime})\in\bar{\mathcal{Y}}\cap C_{D}(0,0)$,
@@
where $\bar{\omega}_{D}$ depends only on $p,d,\lambda,\Lambda,R,D$, $\underset{a\in A}{\sup\,}\left\|g^{a}\right\|_{W^{1,2,p}\left(\mathcal{Y}\cap C_{D+1}(0,0)\right)}$, $\left\|G\right\|_{L^{p}\left(\mathcal{Y}\cap C_{D+1}(0,0)\right)}$, $\left\|u\right\|_{L^{\infty}\left(\mathcal{Y}\cap C_{D+1}(0,0)\right)}$, the modulus of continuity of $b$ in $\partial\mathcal{Y}\cap C_{D+1}(0,0)$, the parameters of the cone condition of $\mathcal{X}$ on $\mathcal{X}\cap B_{D+1}(0)$, and $\left\|u\right\|_{L^{\infty}\left(\mathcal{Y}\cap C_{D+1}(0,0)\right)}$.
@@
Define $\bar{u}:\partial\left(\mathcal{Y}\cap C_{D+1}(0,0)\right)$ as
@@
Similarly, define $\underline{u}:\partial\left(\mathcal{Y}\cap C_{D+1}(0,0)\right)$ as
@@
In particular, note that $\bar{u}$’s and $\underline{u}$’s moduli of continuity are controlled by $\left\|u\right\|_{L^{\infty}\left(\mathcal{Y}\cap C_{D+1}(0,0)\right)}$, and the modulus of continuity of $b$ on $\partial Y\cap C_{D+1}(0,0)$, and that $\bar{u}\geq u\geq\underline{u}$ on $\partial\left(\mathcal{Y}\cap C_{D+1}(0,0)\right)$.
@@
$\begin{cases}&v_{t}+\mathcal{P}^{+}_{\lambda,\Lambda}\left(v_{xx}\right)+R\left|v_{x}\right|\\ &\quad=-\bar{C}\left(1+\left\|u\right\|_{L^{\infty}\left(\mathcal{Y}\cap C_{D+1}(0,0)\right)}+\sup_{a\in A}\left(\left|g^{a}\right|+\left|g^{a}_{t}\right|+\left|g^{a}_{x}\right|+\left|g^{a}_{xx}\right|\right)+G(t,x)\right)\text{ in }\mathcal{Y}\cap C_{D+1}(0,0),\\ &v=\bar{u}\text{ on }\partial\mathcal{Y}\cap C_{D+1}(0,0),\end{cases}$ (11)
@@
$\begin{cases}&v_{t}+\mathcal{P}^{-}_{\lambda,\Lambda}\left(v_{xx}\right)-R\left|v_{x}\right|\\ &\quad=\bar{C}\left(1+\left\|u\right\|_{L^{\infty}\left(\mathcal{Y}\cap C_{D+1}(0,0)\right)}+\sup_{a\in A}\left(\left|g^{a}\right|+\left|g^{a}_{t}\right|+\left|g^{a}_{x}\right|+\left|g^{a}_{xx}\right|\right)+G(t,x)\right)\text{ in }\mathcal{Y}\cap C_{D+1}(0,0),\\ &v=\underline{u}\text{ on }\partial\mathcal{Y}\cap C_{D+1}(0,0).\end{cases}$ (12)
@@
$\left|u(t,x)-b\left(\bar{t},\bar{x}\right)\right|\leq\max\left\{\left|\bar{U}(t,x)-b\left(\bar{t},\bar{x}\right)\right|,\left|\underline{U}(t,x)-b\left(\bar{t},\bar{x}\right)\right|\right\}\leq\bar{\omega}_{boundary}\left(\left|(t,x)-(\bar{t},\bar{x})\right|\right).$ (13)

### v2

@@ 4.2 Stability Theorem
$\begin{cases}u^{n_{j}}_{t}+{F}\left(t,x,u^{n_{j}},u_{x}^{n_{j}},u_{xx}^{n_{j}}\right)\leq-{h^{1}}(t,x)\text{ on }\mathcal{V}_{n_{j}}\\ u^{n_{j}}=g^{1}\text{ on }\partial\mathcal{V}_{n_{j}}.\end{cases}$
By definition of ${h^{1}}$, $g^{1}\in W^{1,2,p}\left(\mathcal{Y}\right)$ is an $L^{p}$-solution of
$\begin{cases}g^{1}_{t}+{F}\left(t,x,g^{1},g^{1}_{x},g^{1}_{xx}\right)=-{h^{1}}(t,x)\text{ on }\mathcal{V}_{n_{j}}\\ g^{1}=g^{1}\text{ on }\partial\mathcal{V}_{n_{j}}.\end{cases}$
Therefore, by our comparison principle (Proposition 1), $g^{1}\leq u^{n_{j}}$ on $\bar{\mathcal{V}}_{n_{j}}$, which yields the desired contradiction. We have thus showed that,for all $j\in\mathbb{N}$, $\mathcal{V}_{n_{j}}=\emptyset$, i.e., $u^{n_{j}}\geq g^{1}$ on $\bar{\mathcal{Y}}$. Taking limits, as $\left(u^{n_{j}}\right)_{j\in\mathbb{N}}$ converges uniformly to $u$ on $\bar{\mathcal{Y}}$, $u\geq g^{1}$ on $\mathcal{Y}$.
We have shown so far that $u=b$ on $\partial\mathcal{Y}$ and that
$\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g^{1}-u\right\}\leq 0\text{ a.e. on }\mathcal{Y},$
i.e., that $u$ is an $L^{p}$-supersolution of (1), with $u=b$ on $\partial\mathcal{Y}$. To conclude, there remains to show that
$\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g^{1}-u\right\}\geq 0\text{ a.e. on }\mathcal{Y}.$
@@
By definition of $\Phi_{\epsilon}$, we have $\Phi_{\epsilon_{n_{j}}}\left(u^{{n_{j}}}(t,x)-g(t,x)\right)\to 1$ pointwise a.e. on $\mathcal{U}$ as $j\to\infty$ Therefore, for all $(t_{0},x_{0})\in\mathcal{U}$ and $D>0$ such that $C_{D}(t_{0},x_{0})\subset\mathcal{U}$, Theorem 2 implies that $u$ is an $L^{p}$-subsolution of
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)=0\text{ on }C_{D}(t_{0},x_{0}).$
@@
$\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g^{1}-u\right\}\geq 0\text{ a.e. on }\mathcal{U}.$
Since $g^{1}\geq u$ on the complement of $\mathcal{U}$, we condlude that
$\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g^{1}-u\right\}\geq 0\text{ a.e. on }\mathcal{Y}.$
and, hence, that $u$ is an $L^{p}$-solution of (1) with $I=1$. This concludes the base case of our induction argument.
Next, suppose that $\bf{P(I)}$ holds for some $I\in\mathbb{N}$. We will show that it also holds for $I+1$. Let $g^{I+1}\in W^{1,2,p}\left(\mathcal{Y}\right)$ with $g^{I+1}\leq b$ on $\partial\mathcal{Y}$. Define $h^{I+1}=-g^{I+1}_{t}-F\left(t,x,g^{I+1},g^{I+1}_{x},g^{I+1}_{xx}\right)$; and consider the following sequence of auxiliary problems. For all $n\in\mathbb{N}$,
$\begin{cases}&\max\bigg{\{}u^{n}_{t}+F\left(t,x,u^{n},u^{n}_{x},u^{n}_{xx}\right)-h^{I+1}(t,x)^{+}\Phi_{\epsilon_{n}}\left(u^{{n}}-g^{I+1}\right)+h^{I+1}(t,x)^{+},\\ &\qquad\qquad\qquad\qquad\qquad\qquad\underset{i\in\{1,\dots,I\}}{\max}\,g^{i}-u^{n}\bigg{\}}=0\text{ on }\mathcal{Y},\\ &u^{n}=b\text{ on }\partial\mathcal{Y}.\end{cases}$ (DI+1)
@@
Proceeding as in the base case $I=1$, we see that the sequence $\left(u^{n}\right)_{n\in\mathbb{N}}$ has a subsequence that converges weakly in $W^{1,2,p}\left(\mathcal{Y}\right)$ and strongly in $C^{0}\left(\bar{\mathcal{Y}}\right)$ to some $u\in W^{1,2,p}\left(\mathcal{Y}\right)$. Moreover, following the same steps as above, one can show that $u$ is an $L^{p}$-solution of (1) with $I+1$ obstacles.
To conclude, there only remains to show that the bound (6) on the $W^{1,2,p}$-norm of $u$ holds. To this end, we first show that $u_{t}=g_{t}$, $u_{x}=g_{x}$, and $u_{xx}=g_{xx}$ a.e. on $\left\{u=g\right\}$. To see this, [^19] A similar argument appears in the proof of Corollary 3.1.2.1 in Evans, (2018). consider $\mathcal{Z}=\left\{(t,x)\in\mathcal{Y}\,:\,u(t,x)=g(t,x)\right\}$. By Lebesgue’s theorem, for almost every $(t,x)\in\mathcal{Z}$,
$\underset{\rho\to 0^{+}}{\lim}\frac{B_{\rho}(t,x)\cap Z}{B_{\rho}(t,x)}=1.$ (8)
Moreover, $u-g\in{W}^{1,2,p}_{loc}(\mathcal{Y})$. Proposition A.1 in Crandall et al., (1998) then implies that, for almost every $(t,x)\in\mathcal{Z}$ and, hence, for a.e. $(t,x)$ such that (8) holds, as $(t^{\prime},x^{\prime})\to(t,x)$,
$(u-g)(t^{\prime},x^{\prime})=(u-g)_{t}(t,x)(\tilde{t}-t)+(u-g)_{x}\cdot(\tilde{x}-x)+o\left(\left|(\tilde{t},\tilde{x})-(t,x)\right|\right).$ (9)
Assume that the vector $\left((u-g)_{t}(t,x),(u-g)_{x}^{\prime}\right)^{\prime}=a\neq 0$ and set
$S=\left\{(v^{t},v^{x})\in\partial B_{1}(0,0)\,:\,a\cdot v\geq\frac{1}{2}\left|a\right|\right\}.$
For each $v\in S$ and $k>0$, set $(\tilde{t},\tilde{x})=(t,x)+kv$ in (9). Then
$(u-g)(t+kv^{t},x+kv^{x})$ $=akv+o(\left|kv\right|)$
$\geq\frac{1}{2}k\left|a\right|+o\left(k\right).$
Therefore, there exists $k_{0}>0$ such that
$(u-g)(t+kv^{t},x+kv^{x})>0,\qquad\forall\,\,0<k<k_{0}\text{ and }v\in S,$
a contradiction to (8). This shows that $a=0$, and $u_{t}(t,x)=g_{t}(t,x)$ and $u_{x}(t,x)=g_{x}(t,x)$ a.e. on $\mathcal{Z}$. By a similar argument, $u_{xx}=g_{xx}$ a.e. on $\left\{(t,x)\in\mathcal{Y}\,:\,u_{x}(t,x)=g_{x}(t,x)\right\}$, hence, a.e. on $\left\{(t,x)\in\mathcal{Y}\,:\,u(t,x)=g(t,x)\right\}$.
Therefore, $u\in{W}^{1,2,p}\left(\mathcal{Y}\right)$ solves
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)=\mathbbm{1}_{\left\{u>\underset{i=1,\dots,I+1}{\max}\,g^{i}(t,x)\right\}}h^{\iota(t,x)}(t,x)-h^{\iota(t,x)}(t,x)$ (10)
where $h^{i}(t,x)=-g^{i}_{t}-F\left(t,x,g^{i},g^{i}_{x},g^{i}_{xx}\right)$ and $\iota(t,x)$ is a measurable selection from $\arg\underset{i\in\{1,\dots I+1\}}{\max}\,g^{i}(t,x)$. The result then follows from Theorem 15.1.3 in Krylov, (2018) (again using Lemma 12.1.9 in Krylov, (2018) to bound $\left\|u\right\|_{L^{\infty}\left(\mathcal{Y}\right)}$).
@@
Suppose that Assumptions 1’, and 27 hold for some $p\in(d+2,\infty)$ and that $\left|A\right|<\infty$. If $u$ is an $L^{p}$-solution of (1), then, for all compact subsets $\mathcal{Y}^{\prime}$ of $\mathcal{Y}$,
@@ 6 Proof of Theorem 1
We first establish our main result for the special case of smooth domains and regular boundary functions $b$ (Section 6.1) and then prove it for the general case (Section 6.2).
6.1 Smooth Domains and Regular Boundary Functions
@@
Suppose that Assumptions 1’, 2’, and 37 hold for some $p\in(d+2,\infty)$. Then, (1) has an $L^{p}$-solution $u$.
@@
The proof of this result builds on Lemma 2. We approximate the obstacle $g=\underset{a\in A}{\sup g^{a}}$ by a sequence of obstacles that satisfy the Assumptions of Lemma 2. We then invoke the stability result for obstacle problems derived in Section 4.2 (Theorem 2) to conclude.
@@
$\begin{cases}\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,\underset{a\in A^{0,n}}{\sup}\,g^{a}-u\right\}=0\text{ on }\mathcal{Y},\\ u=b\text{ on }\partial\mathcal{Y}.\end{cases}$
@@
Since $W^{1,2,p}(\mathcal{Y})$ is separable and reflexive, its closed bounded subsets are weakly sequentially compact by Theorem 1.32 in Demengel et al., (2012). Moreover $W^{1,2,p}\left(\mathcal{Y}\right)$ is compactly embedded in $\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$ by the Rellich-Kondrachov theorem (Theorem 2.84 in Demengel et al., (2012)). Therefore, there exists a function $u\in W^{1,2,p}\left(\mathcal{Y}\right)\cap\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$ and a subsequence $\left(u^{n_{j}}\right)_{j\in\mathbb{N}}\subseteq\left(u^{n}\right)_{n\in\mathbb{N}}$ such that
@@
To conclude, there only remains to show that $u$ solves (1), which follows from Theorem 2. $\blacksquare$
From Corollary 3 and the proof of Lemma 3, we obtain the following interior $W^{1,2,p}$-estimates and $L^{\infty}$ bound.
@@
Suppose that Assumptions 1’, and 28 hold for some $p\in(d+2,\infty)$. If $u$ is an $L^{p}$-solution of (1), then, for all compact subset $\mathcal{Y}^{\prime}$ of $\mathcal{Y}$,
@@
Proof.
By Corollary 2, $u$ is the unique solution of (1). Following the proof of Lemma 3, we note that $u$ is the limit of a sequence of $W^{1,2,p}$ functions, $(u^{n})_{n\in\mathbb{N}}$, such that, for all $n\in\mathbb{N}$, $u^{n}$ solves
$\begin{cases}\max\left\{u^{n}_{t}+F\left(t,x,u^{n},u^{n}_{x},u^{n}_{xx}\right),\,\underset{a\in A^{n}}{\sup}\,g^{a}-u\right\}=0\text{ on }\mathcal{Y},\\ u^{n}=b^{n}\text{ on }\partial\mathcal{Y},\end{cases}$
where $A^{n}$ is finite and the $W^{1,2,p}$ functions $b^{n}$, $n\in\mathbb{N}$, converge to $b$. Corollary 3 then implies that, for all $n\in\mathbb{N}$,
$\left\|u^{n}\right\|_{L^{\infty}\left(\mathcal{Y}\right)}\leq C^{\infty}\left(1+\left\|G\right\|_{L^{p}\left(\mathcal{Y}\right)}+\underset{a\in A}{\sup}\,\left\|g^{a}\right\|_{W^{1,2,p}\left(\mathcal{Y}\right)}+\left\|b\right\|_{L^{\infty}\left(\mathcal{Y}\right)}\right)$
for some $C^{\infty}=C^{\infty}\left(d,p,\lambda,\Lambda,R,T,diam\left(\mathcal{X}\right)\right)\in\mathbb{R}_{+}$. Moreover, for all $n\in\mathbb{N}$ and all compact subset $\mathcal{Y}^{\prime}$ of $\mathcal{Y}$,
$\left\|u^{n}\right\|_{W^{1,2,p}\left(\mathcal{Y^{\prime}}\right)}\leq C\left(1+\underset{a\in A}{\sup}\,\left\|g^{a}\right\|_{W^{1,2,p}\left(\mathcal{Y}\right)}+\left\|G\right\|_{L^{p}\left(\mathcal{Y}\right)}+\frac{1}{dist(\mathcal{Y}^{\prime},\mathcal{Y})}\left\|u\right\|_{L^{\infty}\left(\mathcal{Y}\right)}\right),$
for some $C=C\left(d,p,\lambda,\Lambda,R,diam\left(\mathcal{X}\right),T\right)\in\mathbb{R}_{+}$.
The desired results then follow, since
$\left\|u\right\|_{L^{\infty}\left(\mathcal{Y}\right)}\leq\underset{n\to\infty}{\liminf}\,\left\|u^{n}\right\|_{L^{\infty}\left(\mathcal{Y}\right)}\text{ and }\left\|u\right\|_{W^{1,2,p}\left(\mathcal{Y^{\prime}}\right)}\leq\underset{n\to\infty}{\liminf}\,\left\|u^{n}\right\|_{W^{1,2,p}\left(\mathcal{Y^{\prime}}\right)}.$
@@ 6.2 Proof of Theorem 1
Lemma 3 guarantees that the obstacle problem has an $L^{p}$-solution when $\mathcal{X}$ and $b$ satisfy Assumption 1’ and 2’. To generalize the result to the weaker Assumptions 1 and 2, we study a sequence of equations, each satisfying the assumptions of Lemma 3, that converges to the equation (1). In particular, we approximate (i) $\mathcal{X}$ by a sequence of smooth domains whose cone parameters are uniformly controlled, and (ii) $b$ by a sequence of equicontinuous functions in $W^{1,2,p}\left(\mathcal{Y}\right)$. The $L^{p}$ solutions of the equations in the approximating sequence form an equicontinuous (by Lemma 4, below) and weakly compact (in $W^{1,2,p}_{loc}\left(\mathcal{Y}\right)$) family. The Arzelà-Ascoli theorem then guarantees that a subsequence converges uniformly on the compact subset of $\bar{\mathcal{Y}}$ to some function in $W^{1,2,p}_{loc}\left(\mathcal{Y}\right)\cap\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$. Finally, we invoke the stability result for obstacle problems (Theorem 2) derived in Section 4.2 to conclude.
@@
Suppose that Assumptions 1’, 2’, and 38 hold for some $p\in(d+2,\infty)$. Let $u$ be an $L^{p}$-solution of (1). For all $D>0$, there exists a family of modulus of continuity $\bar{\omega}_{D}$ such that, for all $(t,x),(t^{\prime},x^{\prime})\in\bar{\mathcal{Y}}\cap C_{D}(0,0)$,
@@
where $\bar{\omega}_{D}$ depends only on $p,d,\lambda,\Lambda,R,D$, $\underset{a\in A}{\sup\,}\left\|g^{a}\right\|_{W^{1,2,p}\left(\mathcal{Y}\cap C_{D+1}(0,0)\right)}$, $\left\|G\right\|_{L^{p}\left(\mathcal{Y}\cap C_{D+1}(0,0)\right)}$, $\left\|u\right\|_{L^{\infty}\left(\mathcal{Y}\cap C_{D+1}(0,0)\right)}$, the modulus of continuity of $b$ on $\partial\mathcal{Y}\cap C_{D+1}(0,0)$, and the parameters of the cone condition of $\mathcal{X}$ on $\mathcal{X}\cap B_{D+1}(0)$.
Remark 5
The modulus of continuity $\bar{\omega}_{D}$ for the solution of (1) on $\bar{\mathcal{Y}}\cap C_{D}(0,0)$ given in Lemma 4 only depends on the characteristics of the primitive inside the cylinder $C_{D+1}(0,0)$. It is independent of $\underset{a\in A}{\sup\,}\left\|g^{a}\right\|_{W^{1,2,p}\left(\mathcal{Y}\setminus C_{D+1}(0,0)\right)}$, $\left\|G\right\|_{L^{p}\left(\mathcal{Y}\setminus C_{D+1}(0,0)\right)}$, $\left\|u\right\|_{L^{\infty}\left(\mathcal{Y}\setminus C_{D+1}(0,0)\right)}$, the modulus of continuity of $b$ on $\partial\mathcal{Y}\setminus C_{D+1}(0,0)$, and of the parameters of the cone condition of $\mathcal{X}$ on $\mathcal{X}\setminus B_{D+1}(0)$.
@@
Define $\bar{u}$ on $\partial\left(\mathcal{Y}\cap C_{D+1}(0,0)\right)$ by
@@
Similarly, define $\underline{u}$ on $\partial\left(\mathcal{Y}\cap C_{D+1}(0,0)\right)$ by
@@
We observe that the moduli of continuity of $\bar{u}$ and $\underline{u}$ are controlled by $\left\|u\right\|_{L^{\infty}\left(\mathcal{Y}\cap C_{D+1}(0,0)\right)}$ and the modulus of continuity of $b$ on $\partial Y\cap C_{D+1}(0,0)$, and that $\bar{u}\geq u\geq\underline{u}$ on $\partial\left(\mathcal{Y}\cap C_{D+1}(0,0)\right)$.
@@
$\begin{cases}&v_{t}+\mathcal{P}^{+}_{\lambda,\Lambda}\left(v_{xx}\right)+R\left|v_{x}\right|\\ &\quad=-\bar{C}\left(1+\left\|u\right\|_{L^{\infty}\left(\mathcal{Y}\cap C_{D+1}(0,0)\right)}+\sup_{a\in A}\left(\left|g^{a}\right|+\left|g^{a}_{t}\right|+\left|g^{a}_{x}\right|+\left|g^{a}_{xx}\right|\right)+G(t,x)\right)\text{ on }\mathcal{Y}\cap C_{D+1}(0,0),\\ &v=\bar{u}\text{ on }\partial\mathcal{Y}\cap C_{D+1}(0,0),\end{cases}$ (13)
@@
$\begin{cases}&v_{t}+\mathcal{P}^{-}_{\lambda,\Lambda}\left(v_{xx}\right)-R\left|v_{x}\right|\\ &\quad=\bar{C}\left(1+\left\|u\right\|_{L^{\infty}\left(\mathcal{Y}\cap C_{D+1}(0,0)\right)}+\sup_{a\in A}\left(\left|g^{a}\right|+\left|g^{a}_{t}\right|+\left|g^{a}_{x}\right|+\left|g^{a}_{xx}\right|\right)+G(t,x)\right)\text{ on }\mathcal{Y}\cap C_{D+1}(0,0),\\ &v=\underline{u}\text{ on }\partial\mathcal{Y}\cap C_{D+1}(0,0).\end{cases}$ (14)
@@
$\left|u(t,x)-b\left(\bar{t},\bar{x}\right)\right|\leq\max\left\{\left|\bar{U}(t,x)-b\left(\bar{t},\bar{x}\right)\right|,\left|\underline{U}(t,x)-b\left(\bar{t},\bar{x}\right)\right|\right\}\leq\bar{\omega}_{boundary}\left(\left|(t,x)-(\bar{t},\bar{x})\right|\right)$ (15)